OpenStudy (anonymous):

What is the LCD of 5/6 and -3/8

OpenStudy (anonymous):

i have absoloutley no idea

OpenStudy (anonymous):

i suck at math

OpenStudy (adamaero):

did you look at the examples of your book?

OpenStudy (anonymous):

yes.

OpenStudy (adamaero):

have you tried multiplying both denominators?

OpenStudy (anonymous):

i thought you had to multiply all of the numbers...?

OpenStudy (adamaero):

yes, in a way but anyway, that's the last thing you should do before you ask...

OpenStudy (adamaero):



OpenStudy (anonymous):

so, the LCD is 48?

OpenStudy (adamaero):

8*3 = 48 ?!

OpenStudy (anonymous):

er, i was doing that wrong...

OpenStudy (adamaero):

(8)3 = 8*3 = 8X3

OpenStudy (anonymous):

yeah i was taking 6 and 8 so its 24

OpenStudy (adamaero):

take it in steps my friend and don't take an all or nothing way of thinking. The world is not black and white; you do not 'suck' at math. No one does. We all suck at things we don't know though, and that's why we practice. We try again to learn from mistakes